# Intuitionistic NF set theory

### v5

@@ 5. Order on the cardinals
Proof. By Lemma 5.6, we have $\kappa\leq\mu$. Since $\kappa<\mu$ is defined as $\kappa\leq\mu$ and $\kappa\neq\mu$, it only remains to show $\kappa\neq\mu$. Suppose $\kappa=\mu$. Then $\kappa\leq\lambda$ and $\lambda\leq\kappa$. By Theorem 5.17, we have $\kappa=\lambda$, contradiction. That completes the proof of the lemma.
@@
Proof. Since $\kappa<\lambda$ we have $\kappa\leq\lambda$, by the definition of $<$. Then by Lemma 5.22, $\kappa<\mu$. That completes the proof of the lemma.
@@
Proof. Since $\kappa^{+}\in{\mathbb{F}}$ and $\kappa^{+}$ is inhabited, there exists $x\in\kappa^{+}$. Then $x=y\cup\{c\}$ for some $y\in\kappa$ and $c\not\in x$. Then $x-y=\{c\}$. By Lemma 4.4, since $x\in\kappa^{+}$, $x$ is finite. By Lemma 3.3, $x$ has decidable equality. Therefore $x=y\cup\{c\}=y\cup(x-y)$. Then $y\subseteq x$. It is a proper subset, since $c\in x$ but $c\not\in y$. Now, we will use the right-to-left direction of Lemma 5.8. $\kappa<\kappa^{+}$, substituting $\kappa^{+}$ for $\mu$. That gives us
$\exists x,y\,(x\in\kappa\ \land\ y\in\kappa^{+}\ \land\ x\subset y\ \land\ y=x\cup(y-x)\ {\rightarrow}\ \kappa<\kappa^{+}.$
@@
Since we have verified all four hypotheses, we may conclude $\kappa<\kappa^{+}$. That completes the proof of the lemma.
@@
Proof. Suppose $m\in{\mathbb{F}}$ and $m^{+}\leq m$. By the definition of $\leq$, $m^{+}$ is inhabited. Then by Lemma 5.25, we have $m<m^{+}$, which contradicts Theorem 5.17, since $m^{+}\leq m$. That completes the proof of the lemma.
@@
For $x\in{\mathbb{F}}$, $x\not<x$.
Proof. Immediate from Theorem 5.17, since $x=x$.
@@
Proof. Suppose $x<{\sf zero}$. We will derive a contradiction.
$x\leq{\sf zero}$   by definition of $<$
@@
$b\in\{\Lambda\}$   since ${\sf zero}=\{\Lambda\}$
$b=\Lambda$   by Lemma 2.3
$a=\Lambda$   since $a\subset b$
$\Lambda\in x\cap{\sf zero}$   by definition of intersection
@@
That contradiction completes the proof of the lemma.
@@
Proof. Suppose $\kappa<\mu$. Then there exists $a\in\kappa$ and $b\in\mu$ such that $b=a\cup(b-a)$. Then
@@
$b-a=\Lambda\ \lor\exists u\,(u\in b-a)$   by Lemma 3.4
@@
Case 1, $b-a=\Lambda$. Then $b=a$, so $a\in\kappa\cap\mu$, so by Lemma 5.7, $\kappa=\mu$, contradicting $\kappa<\mu$.
@@
Proof. Suppose $a<b$ and $a,b\in{\mathbb{F}}$. By the definition of $<$, we have $a\leq b$ and $a\neq b$. By the definition of $\leq$, there exists $v\in b$ and $u\in a$ with $u\in SSC(v)$. Then
@@
$v-u\neq\Lambda$   by Lemma 5.7, since $a\neq b$
@@
[…]
Proof. Suppose $\kappa\leq\mu^{+}$. Then by Lemma 5.20,
@@
Proof. Left to right. Suppose $\kappa<\mu^{+}$. Then by the definition of $<$, $\kappa\leq\mu^{+}$ and $\kappa\neq\mu^{+}$. By Lemma 5.31, $\kappa\leq\mu$. By Lemma 5.20, $\kappa<\mu\ \lor\ \kappa=\mu$ as desired.
Right to left. Assume $\mu^{+}\in{\mathbb{F}}$. Then $\mu^{+}$ is inhabited, by Lemma 4.7. If $\kappa=\mu$ then $\kappa<\mu^{+}$ by Lemma 5.25. If $\kappa<\mu$ then $\kappa<\mu^{+}$ by Lemma 5.24. That completes the proof of the lemma.
@@
Proof. By definition, ${\sf zero}=\{\Lambda\}$. Suppose $m\in{\mathbb{F}}$ and $m<{\sf zero}$. By definition of $<$, $m\leq{\sf zero}$ and $m\neq{\sf zero}$. By definition of $\leq$, there exist $a$ and $b$ with $a\in m$ and $b\in{\sf zero}$ and $a\in SSC(b)$. Since ${\sf zero}=\{\Lambda\}$ we have $b=\Lambda$. The only separable subset of $\Lambda$ is $\Lambda$, so $a=\Lambda$. Then by Lemma 5.7, $m={\sf zero}$. But that contradicts $m\neq{\sf zero}$. Therefore the assumptions $m\in{\mathbb{F}}$ and $m<{\sf zero}$ are untenable. That completes the proof of the lemma.
@@
Proof. The formula to be proved is
$\forall x\in{\sf FINITE}\,(x\subseteq{\mathbb{F}}\ {\rightarrow}\ x\neq\Lambda\ {\rightarrow}\ \exists m\in x\,\forall t\,(t\in x\ {\rightarrow}\ t\leq m))$
@@
Base case: immediate, since $\Lambda\neq\Lambda$.
Induction step. Let $x$ be a finite subset of ${\mathbb{F}}$ and $c\in{\mathbb{F}}-x$. By Lemma 3.4, $x$ is empty or inhabited. If $x=\Lambda$, then $c$ is the maximal element of $x\cup\{c\}$, and we are done. So we may assume $x$ is inhabited. Then by the induction hypothesis, $x$ has a maximal element $m$. By Theorem 5.17, $c\leq m$ or $m<c$. If $c\leq m$, then $m$ is the maximal element of $x\cup\{c\}$. If $m<c$, then $c$ is the maximal element of $x\cup\{c\}$, by the transitivity of $\leq$. That completes the proof of the lemma.
@@
Proof. Suppose $x=x^{+}$; then
@@
Now $u\cup\{c\}$ is a finite set, similar to a proper subset of itself (namely $u$). Then by definition, $u\cup\{c\}$ is infinite. By Theorem 3.24, it is not finite. But it is finite. That contradiction completes the proof of the lemma.
@@
Proof. Let $u\in x$ and $u\cup\{c\}\in x^{+}$, with $c\not\in x$. Then by definition of $\leq$ we have $x\leq x^{+}$. By Lemma 5.35, we have $x\neq x^{+}$. Then by definition of $<$, we have $x<x^{+}$. That completes the proof of the lemma.
6. USC, SSC, and similarity
We will replace Rosser and Specker’s use of the full power set $SC$ by the separable power set $SSC$. In this section we prove some lemmas from Specker §2, and some other similar lemmas. For finite sets $a$, since finite sets have decidable equality, every unit subclass is separable, which is helpful. We begin with Specker’s Lemma 2.6, which we take in two steps with the next two lemmas, and after that Specker 2.4 and 2.3.
@@ 6. USC, SSC, and similarity
Let $y\in SSC(USC(a))$. Then there exists $z\in SSC(a)$ such that $y=USC(z)$.
Proof. Suppose $y\in SSC(USC(a)$. Define
(45) $z:=\{u:\{u\}\in y\}.$
That definition is legal since the formula is stratified giving $u$ index 0 and $y$ index 2. Then $y=USC(z)$ since the members of $y$ are the singletons of the members of $z$. I say that $z\subseteq a$: Suppose $u\in z$. Then
@@
$\{u\}\in USC(a)$   since $y\subseteq USC(a)$
$u\in a$   by definition of $USC(a)$
@@
[…]
and that is true since $y$ is a separable subset of $USC(a)$. That completes the proof of the lemma.
@@
$Nc(SSC(USC(a)))=Nc(USC(SSC(a)))$.
Remarks. We follow the proof from [9], p. 368, that Specker cites, checking it constructively with $SSC$ in place of $SC$. But fundamentally, this lemma is just about shuffling brackets. We have $\{\{p\},\{q\},\{r\}\}\in SSC(USC(a))$ corresponding to $\{\{p,q,r\}\}\in USC(SSC(a))$. It is a useful result but not a deep one.
Proof. Let
$W:=\{u:\exists z\,(u=\langle\{z\},USC(z)\rangle)\}.$
The definition is stratified giving $z$ index 1, so $\{z\}$ and $USC(z)$ both get index 2, and $u$ gets index 4. It follows that $W$ is a relation (contains only ordered pairs) and
(46) $\langle x,y\rangle\in W{\ \leftrightarrow\ }\exists z\,(x=\{z\}\ \land\ y=USC(z)).$
I say that $W$ is (the graph of) a one-one-function mapping $USC(SSC(a))$ onto $SSC(USC(a))$. (Formally there is no distinction between a function and its graph.) For if $x$ is given, then $z$ is uniquely determined, so $y$ is uniquely determined; and if $y$ is given with $y=USC(z)$, then $z=\bigcup y$ is unique, so $x=\{z\}$ is unique. Hence $W$ is a function and one-to-one. It remains to show that $W$ is onto. Let $y\in SSC(USC(a))$. By Lemma 6.1, there exists $z\in SSC(a)$ such that $y=USC(z)$. Then $\langle\{z\},y\rangle\in W$. Hence $y$ is in the range of $W$. Since $y$ was an arbitrary member of $SSC(USC(a))$, it follows that $W$ is onto.
We have shown that $W$ is a similarity from $SSC(USC(a))$ to $USC(SSC(a))$. Therefore those two sets have the same cardinal. That completes the proof of the lemma.
@@
Proof. Let $\{a\}$ and $\{b\}$ be unit classes. Define $f=\{\langle a,b\rangle\}$. One can verify that $f:\{a\}\to\{b\}$ is a similarity. We omit the 75 inferences required to do so.
@@
Proof. Let $x\sim\{a\}$. Then let $f:x\to\{a\}$ be a similarity. Since $f$ is onto, there exists $c\in x$ with $f(c)=a$. Let $e\in x$. Then $f(e)\in\{a\}$, so $f(e)=a$. Since $f$ is one to one, $e=c$. Then $x=\{c\}$. That completes the proof of the lemma.
@@
Proof. By definition, ${\sf one}={\sf zero}^{+}$ and ${\sf zero}=\{\Lambda\}$. For any $a$, we have $a\not\in\Lambda$, so
$\Lambda\cup\{a\}=\{a\}\in{\sf zero}^{+}={\sf one}.$
Conversely, if $u\in{\sf one}$, then $u=\Lambda\cup\{a\}$ for some $a$, by definition of successor, so $u=\{a\}$. That completes the proof of the lemma.
@@
$a\in SSC(b)\ {\rightarrow}\ USC(a)\in SSC(SSC(b)).$
Proof. Suppose $a\in SSC(b)$. Since $b$ is finite, it has decidable equality, by Lemma 3.3. Therefore $USC(b)\subseteq SSC(b)$. Since $USC(a)\subseteq USC(b)$, we have
(47) $USC(a)\subseteq SSC(b)$
It remains to show that $USC(a)$ is a separable subset of $SSC(b)$; that is,
$SSC(b)=USC(a)\cup(SSC(b)-USC(a)).$
@@
Right to left: It suffices to show $t\in USC(a)\ {\rightarrow}\ t\in SSC(b)$. Let $t\in USC(a)$. Then $t=\{c\}$ for some $c\in a$. Since $b$ has decidable equality, $t$ is a separable subset of $b$. That completes the right-to-left direction.
Left to right: suppose $t\in SSC(b)$. Then $t\in{\sf FINITE}$, by Lemma 3.19. Then $Nc(t)\in{\mathbb{F}}$, by Lemma 4.21. Then by Lemma 5.18,
$Nc(t)={\sf one}\ \lor\ Nc(t)\neq{\sf one}.$
Case 1, $Nc(t)={\sf one}$. By Lemma 6.5, $t$ is a unit class. Since $a\in SSC(b)$, we have
$x\in b\ {\rightarrow}\ x\in a\ \lor\ x\not\in a.$
Since $t\in USC(a)$ if and only if for some $x$ we have $t=\{x\}\ \land\ x\in a$, we have
$t\in SSC(b)\ {\rightarrow}\ t\in USC(a)\ \lor\ t\not\in USC(a).$
That completes Case 1.
Case 2,$Nc(t)\neq{\sf one}$. Then $Nc(t)$ is not a unit class, by Lemma 6.5 and Lemma 5.7, so the second disjunct on the right holds. That completes the proof of the lemma.
@@
$a\sim b{\ \leftrightarrow\ }USC(a)\sim USC(b).$
Proof. Left-to-right. Suppose $f:a\to b$ is a similarity. Let $g$ be the singleton image of $f$, namely
@@
The definition is legal since the formula is stratified, giving $u$ and $v$ the same index. Then $g:USC(a)\to USC(b)$ is a similarity. We omit the straightforward proof.
Right-to-left. Let $g:USC(a)\to USC(b)$ be a similarity. Define
@@
Again the definition is legal since the formula is stratified, giving $u$ and $v$ the same index. Then $f:a\to b$ is a similarity. We omit the proof.
@@
$a\sim b\ {\rightarrow}\ SSC(a)\sim SSC(b).$
Proof. Let $f:a\to b$ be a similarity. Define
$g:=\{\langle u,f``u\rangle:u\in SSC(a)\}$
where $f``u$ is the image of $u$ under $f$, i.e., the range of the restriction of $f$ to $u$. Then $g:SSC(a)\to SSC(b)$. The fact that the values of $g$ are separable subsets of $b$ follows from Lemma 5.4. We omit the proof that $g$ is one-to-one. To prove $g$ is onto, let $y\in SSC(b)$. Then define
@@
The formula is stratified, giving $u$ and $v$ index 0 and $x$ and $y$ index 1. Hence $x$ can be defined. We omit the proof that $g(x)=y$. ($x$ can also be defined using the operations of domain and inverse relation, which in turn can be defined by stratified comprehension.) That completes the proof (sketch) of the lemma.
@@
If $a$ has decidable equality, then $USC(a)\subseteq SSC(a)$.
Proof. Let $x\in USC(a)$. Then $x=\{u\}$ for some $u\in a$. Then $x\subseteq a$. We must show $a=x\cup(a-x)$. By extensionality, that follows from
@@
which in turn follows from decidable equality on $a$. That completes the proof of the lemma.
@@
$a\subseteq b{\ \leftrightarrow\ }USC(a)\subseteq USC(b),$
Proof. Left to right. Suppose $a\subseteq b$ and $t\in USC(a)$. We must show $t\in USC(b)$. Then $t=\{x\}$ for some $x\in a$. Since $a\subseteq b$ we have $x\in b$. Then $t\in UCS(b)$. That completes the left-to-right direction.
Right to left. Suppose $USC(a)\subseteq USC(b)$ and $t\in a$. We must prove $t\in b$. Since $t\in a$ we have $\{t\}\in USC(a)$. Then $\{t\}\in USC(b)$. Then $\{t\}=\{q\}$ for some $q\in b$. Then $t=q$. Then $t\in b$ as desired. That completes the proof of the lemma.
@@
$a\in SSC(b){\ \leftrightarrow\ }UCS(a)\in SSC(USC(b)).$
Proof. Left to right. Suppose $a\in SSC(b)$. Then $a\subset b$ and
(49) $b$ $=$ $a\cup(b-a).$
@@
(50) $USC(a)\subseteq USC(b)$
It remains to show that $UCS(a)$ is a stable subset of $USC(b)$; that is,
(51) $USC(b)$ $=$ $USC(a)\cup(USC(b)-USC(a)).$

### v6

@@ 5 Order on the cardinals
Proof.
By Lemma 5.6, we have $\kappa\leq\mu$. Since $\kappa<\mu$ is defined as $\kappa\leq\mu$ and $\kappa\neq\mu$, it only remains to show $\kappa\neq\mu$. Suppose $\kappa=\mu$. Then $\kappa\leq\lambda$ and $\lambda\leq\kappa$. By Theorem 5.16, we have $\kappa=\lambda$, contradiction. ∎
@@
Proof.
Since $\kappa<\lambda$ we have $\kappa\leq\lambda$, by the definition of <<. Then by Lemma 5.21, $\kappa<\mu$. ∎
@@
Proof.
Since $\kappa^{+}\in{\mathbb{F}}$ and $\kappa^{+}$ is inhabited, there exists $x\in\kappa^{+}$. Then $x=y\cup\{c\}$ for some $y\in\kappa$ and $c\not\in x$. Then $x-y=\{c\}$. By Lemma 4.4, since $\in\kappa^{+}, xx$ is finite. By Lemma 3.3, $x$ has decidable equality. Therefore $x=y\cup\{c\}=y\cup(x-y)$. Then $y\subseteq x$. It is a proper subset, since $c\in x$ but $c\not\in y$. Now, we will use the right-to-left direction of Lemma 5.7. $\kappa<\kappa^{+}$, substituting $\kappa^{+}$ for $\mu$. That gives us
$\exists x,y\,(x\in\kappa\ \land\ y\in\kappa^{+}\ \land\ x\subset y\ \land\ y=x\cup(y-x)\ {\rightarrow}\ \kappa<\kappa^{+}).$
@@
Since we have verified all four hypotheses, we may conclude $\kappa<\kappa^{+}$. ∎
@@
Proof.
Suppose $m\in{\mathbb{F}}$ and $m^{+}\leq m$. By the definition of $\leq$, $m^{+}$ is inhabited. Then by Lemma 5.26, we have $m<m^{+}$, which contradicts Theorem 5.16, since $m^{+}\leq m$. ∎
@@
For $\in{\mathbb{F}}, xx$$\not<x$.
Proof.
Immediate from Theorem 5.16, since $x=x$. ∎
@@
Proof.
Suppose $x<{\sf zero}$. We will derive a contradiction.
$x\leq{\sf zero}$   by definition of <<
@@
$b\in\{\emptyset\}$   since $\sf zero$$\{\emptyset\}$
$b=\emptyset$   by Lemma 2.3
$a=\emptyset$   since $a\subset b$
$\emptyset\in x\cap{\sf zero}$   by definition of intersection
@@
That is the desired contradiction. ∎
@@
Proof.
Suppose $\kappa<\mu$. Then there exists $a\in\kappa$ and $b\in\mu$ such that $b=a\cup(b-a)$. Then
@@
$b-a=\emptyset\ \lor\exists u\,(u\in b-a)$   by Lemma 3.4
@@
Case 1, $b-a=\emptyset$. Then $b=a$, so $a\in\kappa\cap\mu$, so by Lemma 4.24, $\kappa=\mu$, contradicting $\kappa<\mu$.
@@
Proof.
Suppose $a<b$ and $a,b\in{\mathbb{F}}$. By the definition of <<, we have $a\leq b$ and $a\neq b$. By the definition of $\leq$, there exists $v\in b$ and $u\in a$ with $u\in\mathcal{P}_{s}(v)$. Then
@@
$v-u\neq\emptyset$   by Lemma 4.24, since $a\neq b$
@@
Remark. We cannot replace the \ $\rightarrow$}\ with {\ $\leftrightarrow\$} without the extra assumption, because if $\kappa\leq\mu$ there is no guarantee that $\mu^{+}\in{\mathbb{F}}$.
Proof.
Suppose $\kappa\leq\mu^{+}$. Then by Lemma 5.19,
@@
Proof.
Left to right. Suppose $\kappa<\mu^{+}$. Then by the definition of <<, $\kappa\leq\mu^{+}$ and $\kappa\neq\mu^{+}$. By Lemma 5.32, $\kappa\leq\mu$. By Lemma 5.19, $\kappa<\mu\ \lor\ \kappa=\mu$ as desired.
Right to left. Assume $\mu^{+}\in{\mathbb{F}}$. Then $\mu^{+}$ is inhabited, by Lemma 4.7. If $\kappa=\mu$ then $\kappa<\mu^{+}$ by Lemma 5.26. If $\kappa<\mu$ then $\kappa<\mu^{+}$ by Lemma 5.25. ∎
@@
Proof.
By definition, $\sf zero$$\{\emptyset\}$. Suppose $m\in{\mathbb{F}}$ and $m<{\sf zero}$. By definition of <<, $m\leq{\sf zero}$ and $m\neq{\sf zero}$. By definition of $\leq$, there exist $a$ and $b$ with $a\in m$ and $b\in{\sf zero}$ and $a\in\mathcal{P}_{s}(b)$. Since $\sf zero$$\{\emptyset\}$ we have $b=\emptyset$. The only separable subset of $\emptyset$ is $\emptyset$, so $a=\emptyset$. Then by Lemma 4.24, $m={\sf zero}$. But that contradicts $m\neq{\sf zero}$. Therefore the assumptions $m\in{\mathbb{F}}$ and $m<{\sf zero}$ are untenable. ∎
@@
Proof.
The formula to be proved is
$\forall x\in{\sf FINITE}\,(x\subseteq{\mathbb{F}}\ {\rightarrow}\ x\neq\emptyset\ {\rightarrow}\ \exists m\in x\,\forall t\,(t\in x\ {\rightarrow}\ t\leq m))$
@@
Base case: immediate, since $\emptyset\neq\emptyset$.
Induction step. Let $x$ be a finite subset of $\mathbb{F}$} and $c\in{\mathbb{F}}-x$. By Lemma 3.4, $x$ is empty or inhabited. If $x=\emptyset$, then $c$ is the maximal element of $x\cup\{c\}$, and we are done. So we may assume $x$ is inhabited. Then by the induction hypothesis, $x$ has a maximal element $m$. By Theorem 5.16, $c\leq m$ or $m<c$. If $c\leq m$, then $m$ is the maximal element of $x\cup\{c\}$. If $m<c$, then $c$ is the maximal element of $x\cup\{c\}$, by the transitivity of $\leq$. ∎
@@
Proof.
Suppose $x=x^{+}$; then
@@
Now $u\cup\{c\}$ is a finite set, similar to a proper subset of itself (namely $u$). Then by definition, $u\cup\{c\}$ is infinite. By Theorem 3.24, it is not finite. But it is finite. That contradiction shows $x\neq x^{+}$. ∎
@@
Proof.
Let $u\in x$ and $u\cup\{c\}\in x^{+}$, with $c\not\in x$. Then by definition of $\leq$ we have $x\leq x^{+}$. By Lemma 5.36, we have $x\neq x^{+}$. Then by definition of <<, we have $x<x^{+}$. ∎
6 Power sets and similarity
We will replace Rosser and Specker’s use of the full power set $SC$ by the separable power set $\mathcal{P}_{s}$. In this section we prove some lemmas from Specker §2, and some other similar lemmas. For finite sets $a$, since finite sets have decidable equality, every unit subclass is separable, which is helpful. We begin with Specker’s Lemma 2.6, which we take in two steps with the next two lemmas, and after that Specker 2.4 and 2.3.
@@ 6 Power sets and similarity
Let $y\in\mathcal{P}_{s}(\mathcal{P}_{1}(a))$. Then there exists $z\in\mathcal{P}_{s}(a)$ such that $y=\mathcal{P}_{1}(z)$.
Proof.
Suppose $y\in\mathcal{P}_{s}(\mathcal{P}_{1}(a))$. Define
$z:=\{u:\{u\}\in y\}.$ (46)
That definition is legal since the formula is stratified giving $u$ index 0 and $y$ index 2. Then $y=\mathcal{P}_{1}(z)$ since the members of $y$ are the singletons of the members of $z$. I say that $z\subseteq a$: Suppose $u\in z$. Then
@@
$\{u\}\in\mathcal{P}_{1}(a)$   since $y\subseteq\mathcal{P}_{1}(a)$
$u\in a$   by definition of $\mathcal{P}_{1}(a)$
@@
${\ \leftrightarrow\ }$ $\{u\}\in y\ \lor\{u\}\not\in y$
and that is true since $y$ is a separable subset of $\mathcal{P}_{1}(a)$. ∎
@@
$|\mathcal{P}_{s}(\mathcal{P}_{1}(a))|=|\mathcal{P}_{1}(\mathcal{P}_{s}(a))|$.
Remarks. Or course Specker has $\mathcal{P}$ instead of $\mathcal{P}_{s}$. We follow the proof from [15], p. 368, that Specker cites, checking it constructively with $\mathcal{P}_{s}$ in place of $\mathcal{P}$. But fundamentally, this lemma is just about shuffling brackets. We have $\{\{p\},\{q\},\{r\}\}\in\mathcal{P}_{s}(\mathcal{P}_{1}(a))$ corresponding to $\{\{p,q,r\}\}\in\mathcal{P}_{1}(\mathcal{P}_{s}(a))$. It is a useful result but not a deep one.
Proof.
Let
$W:=\{u:\exists z\,(u=\langle\{z\},\mathcal{P}_{1}(z)\rangle)\}.$
The definition is stratified giving $z$ index 1, so $\{z\}$ and $\mathcal{P}_{1}(z)$ both get index 2, and $u$ gets index 4. It follows that $W$ is a relation (contains only ordered pairs) and
$\langle x,y\rangle\in W{\ \leftrightarrow\ }\exists z\,(x=\{z\}\ \land\ y=\mathcal{P}_{1}(z)).$ (47)
I say that $W$ is (the graph of) a one-one-function mapping $\mathcal{P}_{1}(\mathcal{P}_{s}(a))$ onto $\mathcal{P}_{s}(\mathcal{P}_{1}(a))$. (Formally there is no distinction between a function and its graph.) For if $x$ is given, then $z$ is uniquely determined, so $y$ is uniquely determined; and if $y$ is given with $y=\mathcal{P}_{1}(z)$, then $z=\bigcup y$ is unique, so $x=\{z\}$ is unique. Hence $W$ is a function and one-to-one. It remains to show that $W$ is onto. Let $y\in\mathcal{P}_{s}(\mathcal{P}_{1}(a))$. By Lemma 6.1, there exists $z\in\mathcal{P}_{s}(a)$ such that $y=\mathcal{P}_{1}(z)$. Then $\langle\{z\},y\rangle\in W$. Hence $y$ is in the range of $W$. Since $y$ was an arbitrary member of $\mathcal{P}_{s}(\mathcal{P}_{1}(a))$, it follows that $W$ is onto.
We have shown that $W$ is a similarity from $\mathcal{P}_{s}(\mathcal{P}_{1}(a))$ to $\mathcal{P}_{1}(\mathcal{P}_{s}(a))$. Therefore those two sets have the same cardinal. ∎
@@
Proof.
Let $\{a\}$ and $\{b\}$ be unit classes. Define $f=\{\langle a,b\rangle\}$. One can verify that $f:\{a\}\to\{b\}$ is a similarity. We omit the 75 inferences required to do so. ∎
@@
Proof.
Let $x\sim\{a\}$. Then let $f:x\to\{a\}$ be a similarity. Since $f$ is onto, there exists $c\in x$ with $f(c)=a$. Let $e\in x$. Then $f(e)\in\{a\}$, so $f(e)=a$. Since $f$ is one to one, $e=c$. Then $x=\{c\}$. ∎
@@
Proof.
By definition, ${\sf one}={\sf zero}^{+}$ and $\sf zero$$\{\emptyset\}$. For any $a$, we have $a\not\in\emptyset$, so
$\emptyset\cup\{a\}=\{a\}\in{\sf zero}^{+}={\sf one}.$
Conversely, if $u\in{\sf one}$, then $u=\emptyset\cup\{a\}$ for some $a$, by definition of successor, so $u=\{a\}$. ∎
@@
$a\in\mathcal{P}_{s}(b)\ {\rightarrow}\ \mathcal{P}_{1}(a)\in\mathcal{P}_{s}(\mathcal{P}_{s}(b)).$
Proof.
Suppose $a\in\mathcal{P}_{s}(b)$. Since $b$ is finite, it has decidable equality, by Lemma 3.3. Therefore $\mathcal{P}_{1}(b)\subseteq\mathcal{P}_{s}(b)$. Since $\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{1}(b)$, we have
$\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{s}(b)$ (48)
It remains to show that $\mathcal{P}_{1}(a)$ is a separable subset of $\mathcal{P}_{s}(b)$; that is,
$\mathcal{P}_{s}(b)=\mathcal{P}_{1}(a)\cup(\mathcal{P}_{s}(b)-\mathcal{P}_{1}(a)).$
@@
$t\in\mathcal{P}_{s}(b)$ ${\ \leftrightarrow\ }$ $t\in\mathcal{P}_{1}(a)\ \lor\ (t\in\mathcal{P}_{s}(b)\ \land\ t\not\in\mathcal{P}_{1}(a))$ (49)
Right to left: It suffices to show $t\in\mathcal{P}_{1}(a)\ {\rightarrow}\ t\in\mathcal{P}_{s}(b)$. Let $t\in\mathcal{P}_{1}(a)$. Then $t=\{c\}$ for some $c\in a$. Since $b$ has decidable equality, $t$ is a separable subset of $b$. That completes the right-to-left direction.
Left to right: suppose $t\in\mathcal{P}_{s}(b)$. Then $t\in{\sf FINITE}$, by Lemma 3.19. Then $|t|\in{\mathbb{F}}$, by Lemma 4.21. Then by Lemma 5.17,
$|t|={\sf one}\ \lor\ |t|\neq{\sf one}.$
Case 1, $|t|={\sf one}$. By Lemma 6.5, $t$ is a unit class. Since $a\in\mathcal{P}_{s}(b)$, we have
$x\in b\ {\rightarrow}\ x\in a\ \lor\ x\not\in a.$
Since $t\in\mathcal{P}_{1}(a)$ if and only if for some $x$ we have $t=\{x\}\ \land\ x\in a$, we have
$t\in\mathcal{P}_{s}(b)\ {\rightarrow}\ t\in\mathcal{P}_{1}(a)\ \lor\ t\not\in\mathcal{P}_{1}(a).$
That completes Case 1.
@@
$a\sim b{\ \leftrightarrow\ }\mathcal{P}_{1}(a)\sim\mathcal{P}_{1}(b).$
Proof.
Left-to-right. Suppose $f:a\to b$ is a similarity. Let $g$ be the singleton image of $f$, namely
@@
The definition is legal since the formula is stratified, giving $u$ and $v$ the same index. Then $g:\mathcal{P}_{1}(a)\to\mathcal{P}_{1}(b)$ is a similarity. We omit the straightforward proof.
Right-to-left. Let $g:\mathcal{P}_{1}(a)\to\mathcal{P}_{1}(b)$ be a similarity. Define
@@
Again the definition is legal since the formula is stratified, giving $u$ and $v$ the same index. Then $f:a\to b$ is a similarity. We omit the proof.∎
@@
$a\sim b\ {\rightarrow}\ \mathcal{P}_{s}(a)\sim\mathcal{P}_{s}(b).$
Proof.
Let $f:a\to b$ be a similarity. Define
$g:=\{\langle u,f``u\rangle:u\in\mathcal{P}_{s}(a)\}$
where fuf``u is the image of $u$ under $f$, i.e., the range of the restriction of $f$ to $u$. Then $g:\mathcal{P}_{s}(a)\to\mathcal{P}_{s}(b)$. The fact that the values of $g$ are separable subsets of $b$ follows from Lemma 5.4. We omit the proof that $g$ is one-to-one. To prove $g$ is onto, let $y\in\mathcal{P}_{s}(b)$. Then define
@@
The formula is stratified, giving $u$ and $v$ index 0 and $x$ and $y$ index 1. Hence $x$ can be defined. We omit the proof that $g(x)=y$. ($x$ can also be defined using the operations of domain and inverse relation, which in turn can be defined by stratified comprehension.) ∎
@@
If $a$ has decidable equality, then $\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{s}(a)$.
Proof.
Let $x\in\mathcal{P}_{1}(a)$. Then $x=\{u\}$ for some $u\in a$. Then $x\subseteq a$. We must show $a=x\cup(a-x)$. By extensionality, that follows from
@@
which in turn follows from decidable equality on $a$. ∎
@@
$a\subseteq b{\ \leftrightarrow\ }\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{1}(b),$
Proof.
Left to right. Suppose $a\subseteq b$ and $t\in\mathcal{P}_{1}(a)$. We must show $t\in\mathcal{P}_{1}(b)$. Then $t=\{x\}$ for some $x\in a$. Since $a\subseteq b$ we have $x\in b$. Then $t\in UCS(b)$. That completes the left-to-right direction.
Right to left. Suppose $\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{1}(b)$ and $t\in a$. We must prove $t\in b$. Since $t\in a$ we have $\{t\}\in\mathcal{P}_{1}(a)$. Then $\{t\}\in\mathcal{P}_{1}(b)$. Then $\{t\}=\{q\}$ for some $q\in b$. Then $t=q$. Then $t\in b$ as desired. ∎
@@
$a\in\mathcal{P}_{s}(b){\ \leftrightarrow\ }UCS(a)\in\mathcal{P}_{s}(\mathcal{P}_{1}(b)).$
Proof.
Left to right. Suppose $a\in\mathcal{P}_{s}(b)$. Then $a\subset b$ and
$b$ $=$ $a\cup(b-a).$ (50)
@@
$\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{1}(b)$ (51)
It remains to show that $UCS(a)$ is a stable subset of $\mathcal{P}_{1}(b)$; that is,
$\mathcal{P}_{1}(b)$ $=$ $\mathcal{P}_{1}(a)\cup(\mathcal{P}_{1}(b)-\mathcal{P}_{1}(a)).$ (52)